# HyResPINNs: Adaptive Hybrid Residual Networks for Learning Optimal Combinations of Neural and RBF Components for Physics-Informed Modeling

### v1

@@ Abstract
Physics-informed neural networks (PINNs) are an increasingly popular class of techniques for the numerical solution of partial differential equations (PDEs), where neural networks are trained using loss functions regularized by relevant PDE terms to enforce physical constraints. We present a new class of PINNs called HyResPINNs, which augment traditional PINNs with adaptive hybrid residual blocks that combine the outputs of a standard neural network and a radial basis function (RBF) network. A key feature of our method is the inclusion of adaptive combination parameters within each residual block, which dynamically learn to weigh the contributions of the neural network and RBF network outputs. Additionally, adaptive connections between residual blocks allow for flexible information flow throughout the network. We show that HyResPINNs are more robust to training point locations and neural network architectures than traditional PINNs. Moreover, HyResPINNs offer orders of magnitude greater accuracy than competing methods on certain problems, with only modest increases in training costs. We demonstrate the strengths of our approach on challenging PDEs, including the Allen-Cahn equation and the Darcy-Flow equation. Our results suggest that HyResPINNs effectively bridge the gap between traditional numerical methods and modern machine learning-based solvers.
@@ 1 Introduction
Partial differential equations (PDEs) model a variety of phenomena across science and engineering and are traditionally solved using numerical methods such as finite difference methods (LeVeque, 2007) and finite elements (Strang et al., 1974). Physics-informed neural networks (PINNs) are meshless methods (Raissi et al., 2019; Raissi, 2018) that solve PDEs by training deep feedforward neural networks (DNNs) using PDEs as soft constraints. This traditional PINN training method poses challenges due to complicated loss landscapes arising from the PDE-based soft constraints (Krishnapriyan et al., 2021). Recent approaches to ameliorating issues include curriculum learning (Krishnapriyan et al., 2021), novel optimization techniques (Cyr et al., 2020), domain decomposition (X-PINNs) (Jagtap & Karniadakis, 2020), gradient-enhanced training (G-PINNs) (Yu et al., 2022a), or discretely-trained PINNs using RBF-FD approximations in place of automatic differentiation (DT-PINNs) (Sharma & Shankar, 2022). Further significant strides have been made in developing new DNN architectures to enhance PINNs’ representational capacity, including adaptive activation functions Jagtap et al. (2020b; a), positional embeddings Liu et al. (2020); Wang et al. (2021b), and innovative architectures Wang et al. (2021a); Sitzmann et al. (2020); Gao et al. (2021); Fathony et al. (2020); Moseley et al. (2023); Kang et al. (2023).
In addition to architectural choices, certain training pathologies exist including, spectral bias (Rahaman et al., 2019; Wang et al., 2021b), unbalanced back-propagated gradients (Wang et al., 2021a; 2022b), and causality violations (Wang et al., 2022a; 2024b). Efforts focused on improving PINNs’ training performance include loss re-weighting schemes (Wang et al., 2021a; 2022b; McClenny & Braga-Neto, 2020; 2023; Maddu et al., 2022) and adaptive resampling of collocation points, such as importance sampling (Nabian et al., 2021), evolutionary sampling (Daw et al., ), and residual-based adaptive sampling (Wu et al., 2023).
@@
Combining traditional methods with DNNs is a new line of research which focuses on leveraging the advantages of both. For example, incorporating a Fourier feature layer to preprocess DNN inputs enhances their capacity to capture high-frequency functions (Tancik et al., 2020), while also reducing eigenvector bias in PINNs (Wang et al., 2021b; Raynaud et al., 2022). Fourier neural operators (Li et al., 2020) utilize Fourier layers in the network architecture and have become a popular approach for inverse problems. Chrysos et al. (2022; 2020) augment polynomial neural networks with DNNs, with a direct focus on classification and discriminative problems. Further explorations encompass alternative objective functions, such as those employing numerical differentiation techniques Huang & Alkhalifah (2024) and variational formulations inspired by Finite Element Methods (FEM) Kharazmi et al. (2019); Berrone et al. (2022); Patel et al. (2022), along with additional regularization terms to expedite PINNs’ convergence Yu et al. (2022b); Son et al. (2021).
@@
Recent work has investigated deep architectures for solving PDEs through methods such as stacked DNNs (Howard et al., 2023) and residual blocks with adaptive gating parameters (Wang et al., 2024a). Wang et al. (2024a) incorporate residual blocks with adaptive skip connections, which dynamically balance the input and learned residuals at each layer. Specifically, they incorporate adaptive residual blocks which dynamically balance the contributions of both the input and residual function at each layer, while the stacked approach of  Howard et al. (2023) utilizes transfer learning to improves the ability of PINNs to handle complex, multi-scale PDEs.
While these deep residual based approaches show much promise, the increased architectural complexity, leads to higher computational costs both in terms of memory and training time—requiring careful selection of training routines to prevent instabilities or poor convergence. Finally, while Fourier features help in capturing high-frequency components, they can still struggle with discontinuous solutions or sharp interfaces, which are challenging for neural networks to approximate due to the smoothness of typical activation functions.
@@
To address these challenges, we propose a novel class of network architectures termed HyResPINNs, that combines a standard DNN with a Radial Basis Function (RBF) network within bi-level adaptive residual blocks, such that the relative contribution of both the DNN and RBF networks are adaptively learned during training, along with the full residual block outputs. The proposed architecture leverages the strengths of both smooth and non-smooth function approximators to model complex physical systems. The standard DNN using smooth activation functions effectively captures the continuous, global behaviors of the solution, which are common in many physical phenomena. While the RBF network then captures the localized, discontinuous or sharp features in the target solution. This representational division ensures that the model can accurately capture both smooth and non-smooth components of the solution, forming a robust network capable of modeling many types of physical systems.
Our main contributions are summarized as follows:
@@
  • Propose a novel adaptive residual architecture: We introduce a new class of physics-informed neural networks, HyResPINN, that combine standard neural networks with Radial Basis Function (RBF) networks within bi-level adaptive residual blocks.
  • Demonstrate the superiority of the block structure: We show that our residual block architecture provides significant improvements over standard approaches in capturing both smooth and non-smooth features, leading to more accurate modeling of complex physical systems.
  • Highlight the benefits of adaptivity between residual blocks: We demonstrate that the adaptive learning of contributions between the NN and RBF networks, as well as between residual blocks, results in superior performance and stability compared to non-adaptive methods.
  • Thorough empirical evaluation: Show our method HyResPINN  outperforms standard PINNs along with state of the art methods in PINNs on an array of baseline problems, confirming the general applicability of HyResPINNs.
@@ 2 Background
2.1 Physics-informed neural network
Given the spatio-temporal domain $\Omega\subset\mathbb{R}^{d}$ defined on $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$ where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with regular enough boundary $\partial\Omega$, the general form of a parabolic PDE is,
$\mathbf{u}_{t}+{\mathcal{F}}[\mathbf{u}]=\mathbf{f}\,,$ (1)
such that ${\mathcal{F}}[\cdot]$ is a linear or nonlinear differential operator, and $\mathbf{u}(t,\mathbf{x})$ denotes a unknown solution.
The general initial and boundary conditions can be then formulated as:
$\mathbf{u}(0,\mathbf{x})=\mathbf{g}(\mathbf{x}),\quad\mathbf{x}\in\Omega\,,$ (2)
$\mathcal{B}[\mathbf{u}]=0,\quad t\in[0,T],\ \mathbf{x}\in\partial\Omega\,.$ (3)
Here, $\mathbf{f}$ and $\mathbf{g}(\mathbf{x})$ are given functions with certain regularity; $\mathcal{B}[\cdot]$ denotes an abstract boundary operator, representing various boundary conditions such as Dirichlet, Neumann, Robin, and periodic conditions.
We aim at approximating the unknown solution $\mathbf{u}(t,\mathbf{x})$ by a deep neural network $\mathbf{u}_{\mathbf{\theta}}(t,\mathbf{x})$, where $\mathbf{\theta}$ denotes the set of all trainable parameters of the network (e.g., weights and biases). If a smooth activation function is employed, $\mathbf{u}_{\mathbf{\theta}}$ provides a smooth representation that can be queried for any $(t,\mathbf{x})$. The PDE residuals are defined as,
$\mathcal{R}_{{\rm int}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial\mathbf{u}_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{D}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})-\mathbf{f}(\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\Omega\,,$ (4)
and spatial and temporal boundary residuals, respectively, by
$\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t,\mathbf{x})=\mathcal{B}[\mathbf{u}_{\theta}](t,\mathbf{x})\,,\quad(t,\mathbf{x})\in[0,T]\times\partial\Omega\,,$ (5)
and
$\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x})=\mathbf{u}_{\theta}(0,\mathbf{x})-\mathbf{g}(\mathbf{x})\,,\quad\mathbf{x}\in\Omega\,.$ (6)
@@ 2.1 Physics-informed neural network
$\mathcal{L}(\mathbf{\theta}):=\underbrace{\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[\mathbf{u}_{\mathbf{\theta}}](\mathbf{x}_{ic}^{i})\right|^{2}}_{\mathcal{L}_{ic}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left|\mathcal{R}_{{\rm bc}}[\mathbf{u}_{\mathbf{\theta}}](t_{bc}^{i},\mathbf{x}_{bc}^{i})\right|^{2}}_{\mathcal{L}_{bc}(\mathbf{\theta})}+\underbrace{\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{\rm int}[\mathbf{u}_{\mathbf{\theta}}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}}_{\mathcal{L}_{r}(\mathbf{\theta})}\,,$ (7)
which aims to enforce the neural network function $\mathbf{u}_{\theta}$ to satisfy the PDEs (1) with initial and spatial boundary conditions (2)–(3). The training data points $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be the vertices of a fixed mesh or points randomly sampled at each iteration of a gradient descent algorithm.
2.2 Radial Basis Function Networks
The general form of an RBF network for the training data points $\mathbf{x}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with corresponding function values $y_{i}$ is given by,
$f({\bf x})$ $=\sum_{j=1}^{N_{c}}c_{j}\phi(||{\bf x}-{\bf x}_{j}^{c}||),$ (8)
where ${\bf x}^{c}=\{{\bf x}_{j}^{c}\}_{j=1}^{N_{c}}$ are the center points of the radial basis functions (RBFs). Here, $\phi$ represents the radial basis function, which are commonly chosen as the Gaussian, multi-quadric, or inverse multi-quadric functions, each offering different characteristics and benefits. The elements of the kernel matrix are given by:
$\mathbf{K}_{ij}$ $=\phi(||{\bf x}_{i}-{\bf x}^{c}_{j}||),$ (9)
where ${\bf x}_{i}$ represent the data points, and (9) evaluates the RBF based on the distance between the points. The kernel matrix captures the pairwise interactions between data points through the chosen RBF, and its structure is typically symmetric and positive-definite, assuming an appropriate choice of $\phi$. The choice of center points can vary depending on the application; in some cases, the center points may coincide with the data points, but they may also be selected independently of the data points to optimize approximation accuracy or computational efficiency. The RBFs determine how much influence each center has on the input, based on the distance between the input and the center.
@@ 2.2 Radial Basis Function Networks
Figure 1: Illustration of the RBF+PINN hybrid residual block with trainable strength connections between the RBF and PINN outputs.
Gradient descent is sometimes used to optimize both the centers and the unknown weights ($c_{j}$) simultaneously. These approaches are referred to as RBF networks, where the radial basis functions are the hidden layer activations, and the RBF coefficients are the trainable network parameters. The mean-squared error between the linear combination of the hidden layer outputs and the known functions values is commonly minimized.
@@ 3 Hybrid Residual PINNs (HyResPINNs)
In this section, we describe our proposed architecture—HyResPINNs. HyResPINNs  are a novel type of residual network such that each residual block incorporates an RBF kernel along with a standard DNN.
3.1 Hybrid Residual Blocks
Figure 1 shows the hybrid residual block specifics. We denote the output of the $l$-th residual block for the input ${\bf x}^{(l)}$ as $H^{(l)}({\bf x}^{(l)})$. Formally, the forward pass of each hybrid residual block in the HyResPINN  architecture is,
$H^{(l)}({\bf x}^{(l)})=\sigma\big{(}\phi(\alpha^{(l)})F_{R}^{(l)}({\bf x}^{(l)})+(1-\phi(\alpha^{(l)}))F_{N}^{(l)}({\bf x}^{(l)})\big{)},$ (10)
such that $\sigma$ is some non-linear activation function, $\alpha^{(l)}\in\mathbb{R}$ is a trainable parameter, $\phi$ represents the sigmoid function, $F_{R}^{(l)}({\bf x}^{(l)})$ is the output of the RBF network and $F_{N}^{(l)}({\bf x}^{(l)})$ is the output from a DNN or PINN within the $l$-th block.
[…]
@@ 3.1 Hybrid Residual Blocks
Figure 2: Illustration of the HyResPINN  architecture using three blocks.
Adaptive Hybrid Contribution Coefficients
The trainable parameter $\alpha^{(l)}$ controls the relative contributions of the DNN and RBF components to the output—where optimizing $\alpha$ finds the best balance between these two elements. Since the sigmoid function $\phi$ constrains its output to the range $[0,1]$, the resulting output is a convex combination of the RBF and DNN components. We further incorporate adaptable residual connection parameters between each hybrid block, denoted as $\beta^{(l)}$—similar to the approaches in (Howard et al., 2023) and (Wang et al., 2024a)
However, unlike (Wang et al., 2024a), where their adaptable parameters are initialized to zero to force the network to learn non-linearities from scratch, we initialize each $\alpha^{(l)}=0.5$, ensuring equal contributions from both components at the start of training, and each $\beta^{(l)}=1$. When $\alpha=1$, only the RBF network contributes; when $\alpha=0$, only the DNN contributes. The optimal choice of $\alpha$ depends on the problem characteristics. For problems with large regions of smoothness, the model might favor a lower $\alpha$, assigning more weight to the DNN components in each residual block. Conversely, for problems involving sharp transitions or discontinuities, a higher $\alpha$ (favoring the RBF) may be preferable. In hybrid problems with smooth and non-smooth regions, $\alpha$ will likely fall between 0.4 and 0.6, providing a balanced combination of the two networks.
To encourage smoother solutions and prevent the network from introducing excessive non-linearity, we add a regularization term to each block’s trainable parameter $\alpha$. This regularization penalizes large values of $\alpha$, effectively controlling the contribution of non-linearity from the RBF components. Specifically, the total loss function is defined as:
$\mathcal{L}=\lambda_{ic}\mathcal{L}_{ic}({\boldsymbol{\theta}})+\lambda_{bc}\mathcal{L}_{bc}({\boldsymbol{\theta}})+\lambda_{r}\mathcal{L}_{r}({\boldsymbol{\theta}})+\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2},$ (11)
where $\mathcal{L}_{ic}$ and $\mathcal{L}_{bc}$ represent the loss terms for the initial and boundary conditions, $\mathcal{L}_{r}$ represents the residual loss, and the final term $\lambda_{p}\sum_{i=1}^{N_{blocks}}\alpha_{i}^{2}$ acts as an $L_{2}$ regularization on the $\alpha$ parameters. The regularization helps balance the contributions of the smooth neural network and non-linear RBF components, promoting smoother solutions and more stable training.
Figure 3: 2D Allen-Cahn equation: Comparison of (a subset of) the learned RBF kernels for a standard RBF network (top row) and the proposed hybrid RBF+NN residual block approach (bottom row). Each subplot shows the RBF value corresponding to different learned RBF centers in the input domain, marked by red crosses.
Adaptive RBF Block Kernel
The output of the RBF kernel $F_{R}^{(l)}({\bf x}^{(l)})$ can be described as,
$F_{R}^{(l)}({\bf x}^{(l)})=K({\bf x})W,$ (12)
such that the kernel matrix $K$ is defined as (9) where the kernel matrix $K$ is based on the Wendland $C^{4}$ kernel in 2D. In this work, we focus on isotropic version of the Wendland kernel.
The isotropic Wendland $C^{4}$ kernel for input ${\bf x}$ and the $i$-th center is defined as:
$\phi_{i}(\mathbf{x})=\left(1-\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)_{+}^{6}\left(35\left(\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}\right)^{2}+18\frac{\|{\bf x}-{\bf x}^{c}_{i}\|}{\tau_{i}}+3\right),$ (13)

### v2

# HyResPINNs: Hybrid Residual Networks for Adaptive Neural and RBF Integration in Solving PDEs

@@ Abstract
Physics-informed neural networks (PINNs) have emerged as a powerful approach for solving partial differential equations (PDEs) by training neural networks with loss functions that incorporate physical constraints. In this work, we introduce HyResPINNs, a novel class of PINNs featuring adaptive hybrid residual blocks that integrate standard neural networks and radial basis function (RBF) networks. A distinguishing characteristic of HyResPINNs is the use of adaptive combination parameters within each residual block, enabling dynamic weighting of the neural and RBF network contributions. Our empirical evaluation of a diverse set of challenging PDE problems demonstrates that HyResPINNs consistently achieve superior accuracy to baseline methods. These results highlight the potential of HyResPINNs to bridge the gap between classical numerical methods and modern machine learning-based solvers, paving the way for more robust and adaptive approaches to physics-informed modeling.
@@ 1 Introduction
Partial differential equations (PDEs) model a variety of phenomena across science and engineering and are traditionally solved using numerical methods such as finite difference methods (LeVeque, 2007) and finite elements (Strang et al., 1974). Physics-informed neural networks (PINNs) are meshless methods (Raissi et al., 2019b; Raissi, 2018) that solve PDEs by training deep feedforward neural networks (DNNs) using PDEs as soft constraints. This traditional PINN training method poses challenges due to complicated loss landscapes arising from the PDE-based soft constraints (Krishnapriyan et al., 2021). Recent approaches to ameliorating issues include curriculum learning (Krishnapriyan et al., 2021), novel optimization techniques (Cyr et al., 2020), domain decomposition (X-PINNs) (Jagtap & Karniadakis, 2020), gradient-enhanced training (G-PINNs) (Yu et al., 2022a), or discretely-trained PINNs using RBF-FD approximations in place of automatic differentiation (DT-PINNs) (Sharma & Shankar, 2022). Further significant strides have been made in developing new DNN architectures to enhance PINNs’ representational capacity, including adaptive activation functions Jagtap et al. (2020b; a), positional embeddings Liu et al. (2020); Wang et al. (2021b), and innovative architectures Wang et al. (2021a); Sitzmann et al. (2020); Gao et al. (2021); Fathony et al. (2020); Moseley et al. (2023); Kang et al. (2023).
@@
Combining traditional methods with DNNs is a growing area of research that seeks to leverage the advantages of both approaches. For instance, Fourier feature layers have been incorporated into DNNs to improve their ability to represent high-frequency functions (Tancik et al., 2020) while also reducing eigenvector bias in PINNs (Wang et al., 2021b; Raynaud et al., 2022). Fourier neural operators (Li et al., 2020) extend this idea by incorporating Fourier layers within network architectures, which have proven useful for solving inverse problems. Similarly, physics-informed radial basis networks (PIRBNs) have been developed as efficient local approximators that integrate domain knowledge, particularly excelling in solving PDEs with high-frequency features or ill-posed domains (Bai et al., 2023; Fu et al., 2024). Other works have explored polynomial-based neural networks (Chrysos et al., 2022; 2020) and alternative optimization techniques, such as numerical differentiation methods Huang & Alkhalifah (2024) and variational formulations inspired by Finite Element Methods (FEM) Kharazmi et al. (2019); Berrone et al. (2022); Patel et al. (2022). Additional regularization strategies have also been proposed to improve the convergence behavior of PINNs Yu et al. (2022b); Son et al. (2021).
@@
Recent studies have investigated the use of deep architectures for solving PDEs through methods such as stacked DNNs (Howard et al., 2023) and residual networks with adaptive gating parameters (Wang et al., 2024) (PirateNets). PirateNets introduce adaptive residual blocks that dynamically balance the contributions of input features and learned residuals at each layer. Meanwhile, stacked PINNs utilize transfer learning to enhance PINNs’ ability to solve complex, multi-scale PDEs.
Despite these advances, several challenges remain. Residual-based deep architectures often introduce significant computational costs due to increased memory and training time requirements. These approaches require careful tuning of training routines to prevent instabilities or slow convergence. Additionally, while Fourier-based methods improve the ability of DNNs to capture high-frequency components, they still struggle with discontinuous solutions or sharp interfaces—structures that neural networks inherently struggle to approximate due to the smoothness constraints imposed by typical activation functions. Moreover, approaches such as PirateNets rely on residual connections initialized to zero, forcing the network to learn complex function structures from scratch rather than leveraging prior knowledge from earlier layers. Since these methods rely entirely on DNN-based architectures, they lack explicit mechanisms for handling both smooth and non-smooth function components in a structured way.
@@
To address these challenges, we propose a novel class of network architectures termed HyResPINNs, which introduces a hybrid function approximation framework by integrating standard DNNs with Radial Basis Function (RBF) networks within bi-level adaptive residual blocks. Unlike existing residual architectures, HyResPINN adaptively learns both the relative contributions of the DNN and RBF components within each block, as well as the full residual block outputs throughout training. This hybrid formulation enables the model to leverage the strengths of both smooth and non-smooth function approximators to more effectively capture complex physical systems. The standard DNN, utilizing smooth activation functions, is well-suited for learning continuous, global behaviors commonly observed in many physical phenomena. Meanwhile, the RBF network component provides localized function approximation, effectively capturing sharp interfaces, discontinuities, and high-gradient regions that are difficult for purely neural network-based approaches to represent. By combining these two complementary representations in an adaptive manner, the proposed architecture ensures that the model can simultaneously handle both smooth and non-smooth components of the solution. This results in a more robust and expressive model for solving PDEs across a wide range of problem domains.
We highlight the following key contributions:
@@
  • A novel hybrid residual architecture: We introduce HyResPINNs, a new class of PINNs that combine deep neural networks with radial basis function networks, enabling an adaptive balance between global smoothness and localized feature representations.
  • Bi-level adaptivity through convex combination residual learning: Unlike prior residual-based approaches that use fixed or additive skip connections, we propose a convex combination formulation that dynamically learns the optimal contribution of each residual block output and its input, resulting in improved stability and adaptability during training.
  • Robust handling of both smooth and non-smooth features: The proposed architecture effectively captures both globally smooth behaviors and sharp, localized structures by leveraging the complementary strengths of neural networks and RBFs, outperforming existing methods in modeling multi-scale and discontinuous solutions.
  • Improved training stability through regularization: We introduce an explicit regularization mechanism on the adaptive gating parameters, preventing excessive dominance of either the DNN or RBF components, thereby ensuring a well-balanced hybrid function approximation.
  • Comprehensive empirical evaluation: We demonstrate the effectiveness of HyResPINNs  across a range of benchmark PDE problems, showing superior accuracy, efficiency, and stability compared to standard PINNs and state-of-the-art baseline architectures.
Through these contributions, we establish HyResPINNs  as a principled and effective approach for learning solutions to PDEs, providing a flexible framework that generalizes well across different problem domains.
@@ 2 Background
Physics-Informed Neural Networks.
Given the spatio-temporal domain $\Omega\subset\mathbb{R}^{d}$ with spatial dimension $d$, defined on $[0,T]\times\Omega\subset\mathbb{R}^{1+d}$ where $\Omega$ is a bounded domain in $\mathbb{R}^{d}$ with boundary $\partial\Omega$, the general form of a parabolic PDE is, $u_{t}+{\mathcal{F}}[u]=f$, such that ${\mathcal{F}}[\cdot]$ is a linear or nonlinear differential operator, and $u(t,\mathbf{x})$ denotes a unknown solution. The general initial conditions are $u(0,\mathbf{x})=g(\mathbf{x})$, for $\mathbf{x}\in\Omega$, a boundary conditions $\mathcal{B}[\mathbf{u}]=0$, for $\quad t\in[0,T]$ and $\mathbf{x}\in\partial\Omega$. Here, $f$ and $g(\mathbf{x})$ are given and $\mathcal{B}[\cdot]$ denotes an abstract boundary operator.
The general formulation of PINNs Raissi et al. (2019a) aim to approximate the unknown solution $u(t,\mathbf{x})$ by a deep neural network $u_{\theta}(t,\mathbf{x})$, where $\theta$ denotes the set of all trainable parameters of the network. The PDE residuals are defined as,
$\mathcal{R}_{r}[u_{\mathbf{\theta}}](t,\mathbf{x})=\frac{\partial u_{\mathbf{\theta}}}{\partial t}(t,\mathbf{x})+\mathcal{F}[u_{\mathbf{\theta}}](t,\mathbf{x})-f(\mathbf{x})\,.$ (1)
The boundary and initial condition residuals are,
$\mathcal{R}_{b}[u_{\theta}](t,\mathbf{x})$ $=\mathcal{B}[u_{\theta}](t,\mathbf{x})\,,$ (2)
$\mathcal{R}_{i}[u_{\theta}](\mathbf{x})$ $=u_{\theta}(0,\mathbf{x})-g(\mathbf{x})\,.$ (3)
@@
$\mathcal{L}(\theta):=\frac{\lambda_{ic}}{N_{ic}}\sum_{i=1}^{N_{ic}}\left|\mathcal{R}_{{\rm ic}}[u_{\theta}](\mathbf{x}_{ic}^{i})\right|^{2}$ (4)
$+\frac{\lambda_{b}}{N_{b}}\sum_{i=1}^{N_{b}}\left|\mathcal{R}_{b}[u_{\theta}](t_{bc}^{i},\mathbf{x}_{b}^{i})\right|^{2}$ $+\frac{\lambda_{r}}{N_{r}}\sum_{i=1}^{N_{r}}\left|\mathcal{R}_{r}[u_{\theta}](t^{i}_{r},\mathbf{x}^{i}_{r})\right|^{2}\,,$ (5)
which aims to enforce the neural network function $u_{\theta}$ to satisfy the PDE with initial and boundary conditions. $\lambda_{r}$, $\lambda_{b}$, and $\lambda_{i}$ are typically hyperparameters that control the interplay between the different loss terms. The training data points $\{\mathbf{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$, $\{t_{bc}^{i},\mathbf{x}_{bc}^{i}\}_{i=1}^{N_{bc}}$ and $\{t_{r}^{i},\mathbf{x}_{r}^{i}\}_{i=1}^{N_{r}}$ can be fixed mesh or points randomly sampled at each iteration of a gradient descent algorithm.
Radial Basis Function Networks.
The general form of a radial basis function (RBF) network for a set of training data points $\mathbf{x}=\{\mathbf{x}_{i}\}_{i=1}^{N}$ with corresponding function values $y_{i}$ is given by
$f(\mathbf{x})$ $=\sum_{j=1}^{N_{c}}c_{j}\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||),$ (6)
where $\mathbf{x}^{c}=\{\mathbf{x}_{j}^{c}\}_{j=1}^{N_{c}}$ represents the center points of the radial basis functions (RBFs), and $\psi$ is the chosen radial basis function. Common choices for $\psi$ include the Gaussian, multi-quadric, and inverse multi-quadric functions, each offering different approximation properties.
The kernel matrix $\mathbf{K}\in\mathbb{R}^{N\times N_{c}}$ is constructed as $K_{ij}=\psi(||\mathbf{x}_{i}-\mathbf{x}_{j}^{c}||)$, where $\mathbf{x}_{i}$ represents the data points, and $K_{ij}$ evaluates the RBF response based on the distance between input points and center locations. The kernel matrix captures pairwise interactions between data points and is typically symmetric and positive-definite, assuming an appropriate choice of $\psi$. The selection of center points $\mathbf{x}^{c}$ plays a crucial role in the model’s performance. The RBFs determine the influence of each center on the input space based on the distance between $\mathbf{x}$ and $\mathbf{x}_{j}^{c}$.
Gradient-based optimization is sometimes used to update both the center locations $\mathbf{x}^{c}$ and the RBF coefficients $c_{j}$ during training. This adaptive learning process defines an RBF network, where the radial basis functions act as the hidden layer activations, and the RBF coefficients serve as the trainable network parameters. The network is typically trained by minimizing the mean squared error (MSE) between the predicted function values and the given training data.
@@
Figure 1: Illustration of the hybrid residual block with trainable strength connections between the RBF and NN outputs.
@@ 3 Hybrid Residual PINNs (HyResPINNs)
We propose HyResPINN, a deep learning framework that integrates radial basis function networks (RBFs) and feedforward neural networks (NNs) within a hybrid residual structure. This architecture enhances function approximation by leveraging localized adaptivity from RBFs and global smoothness from NNs.
Hybrid Block $1$Hybrid Block $2$ Hybrid Block $3$ $\mathbf{x}$$\bar{\mathbf{a}}^{(1)}$$\bar{\mathbf{a}}^{(2)}$$\bar{\mathbf{a}}^{(3)}$$\mathbf{y}$$\phi(\beta^{(1)})$$\phi(\beta^{(2)})$$\phi(\beta^{(3)})$
@@
Figure 2: Illustration of the HyResPINN architecture using three hybrid residual blocks. The trainable skip connections $\phi(\beta^{(l)})$ modulate the information flow between blocks.
3.1 Hybrid Residual Block Structure
The core component of HyResPINN is the hybrid residual block, illustrated in Figure 1. Each block consists of two primary components: an RBF-Net, denoted as $F_{R}$, which captures localized behavior via an RBF kernel matrix, and an NN component, denoted as $F_{N}$, which provides a globally smooth representation of the solution. These two outputs are combined using a trainable weighting function $\phi(\alpha^{(l)})$, which dynamically adjusts the contributions of the RBF-Net and NN within each block.
Given a set of $N$ points, $\mathbf{x}\in\mathbb{R}^{N\times d}$, the output of the $l$-th hybrid block is computed as:
$F^{(l)}(\mathbf{x})=\phi(\alpha^{(l)})F_{R}^{(l)}(\mathbf{x})+\left(1-\phi(\alpha^{(l)})\right)F_{N}^{(l)}(\mathbf{x}),$ (7)
where $F_{R}^{(l)}(\mathbf{x})\in\mathbb{R}^{N\times p}$ represents the output of the RBF-Net, while $F_{N}^{(l)}(\mathbf{x})\in\mathbb{R}^{N\times p}$ corresponds to the NN output. Similar to Howard et al. (2023), the function $\phi(\alpha^{(l)})$ is a learnable gating mechanism that determines the relative contributions of the two models. To introduce non-linearity between block outputs and constrain them within a well-defined range, each block output $F^{(l)}$ is passed through a nonlinear activation function $\sigma(\cdot)$. In this work, $\sigma=\tanh$ is chosen to ensure that the outputs remain within the bounded space $[-1,1]^{d}$.
RBF Block Component.
The RBF-Net component is defined by the function:
$F_{R}^{(l)}(\mathbf{x})=W^{(l)}\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||),$ (8)
where $W^{(l)}\in\mathbb{R}^{N\times p}$ are the trainable RBF coefficients, $\mathbf{x}_{j}^{c}$ represents the RBF centers, and $\psi(||\mathbf{x}-\mathbf{x}_{j}^{c}||)$ is a radial basis function that applies a distance-based transformation to the input $\mathbf{x}$. The output dimension of the block is denoted by $p$, which serves as a user-defined hyperparameter. This component provides localized adaptivity, allowing the network to capture fine details and sharp transitions in the function approximation. The choice of $\psi$ influences the behavior of the model, with common options including Gaussian, multi-quadric, and inverse multi-quadric functions.
Wendland Kernel Selection.
In this work, we focus on an isotropic Wendland $C^{4}$ kernel, which is compactly supported, ensuring both sparsity and computational efficiency. The Wendland kernel is defined as:
$\psi_{i}(\mathbf{x})$ $=\left(1-\frac{r_{i}}{\tau_{i}}\right)_{+}^{6}\left(35\left(\frac{r_{i}}{\tau_{i}}\right)^{2}+18\frac{r_{i}}{\tau_{i}}+3\right),$ (9)